OpenStudy (anonymous):

Write a quadratic equation having the given numbers as solutions. -5 and -2

jimthompson5910 (jim_thompson5910):

x=-5 or x=-2 x-(-5)=0 or x-(-2)=0 x+5=0 or x+2=0 (x+5)(x+2)=0 x^2+2x+5x+10=0 x^2+7x+10=0

OpenStudy (anonymous):

(x+5)(x+2) = 0 has -5 and -2 as solutions right? Now just FOIL that out.

OpenStudy (anonymous):

\[(x+5)(x+2)\] \[x ^{2}+7x+10\]